# Locate-then-edit for Multi-hop Factual Recall under Knowledge Editing

### v2

@@ 17       $W^{l}\leftarrow W^{l}+\Delta^{l}$
where $K_{0}=\left[k^{1}_{0}\,|\,k^{2}_{0}\,|\,\cdots\,|\,k^{N}_{0}\right]$ and $V_{0}=W^{out}K_{0}$ contain all the knowledge we want to preserve, $K_{E}=\left[k^{1}_{e}\,|\,k^{2}_{e}\,|\,\cdots\,|\,k^{E}_{e}\right]$ is the matrix containing the edits we try to make and $V_{e}=[v^{*}_{e_{1}}\ |\dots\ |\ v^{*}_{e_{E}}]$ represents the target representations of the new knowledge. ($K_{E},V_{E}$) corresponds to the edited fact set $\{\left(s_{i},r_{i},o^{*}_{i}\right)|i=1,2,\cdots,E\}$. We consider the target weight $\hat{W}$ as the sum of the original weight $W^{out}$ and the incremental weight $\Delta$, as explicated in Li et al. (2024c), a closed-form solution to the incremental weight can be derived:
@@
The primary differences between the first and furtherance edits are reflected in the edit prompt $T_{e}$ and the layers edited. For example, for the edit instance $e=(s,r,o\to o*)$, the first edit utilized a one-hop edit template $T_{e}=T_{r}(s)$ provided by MQuAKE to edit layers [3,8] of the GPT-J model in the subject last token position. For the furtherance edit, a two-hop template $T_{e}=T_{C}(s^{\prime})$ composed of a support case $(s^{\prime},r,s)$ and $(s,r,o^{*})$, and this two-hop template was applied to edit layers [16,20] of the GPT-J model in the last token position.
Appendix D Addition Experimental Settings
D.1 Criteria for classifying dataset into pre, mid, and post.
Consider a multi-hop question composed of $n$ triples. We define the positions of edits (with index starting from 1) as the set $\{e_{1},e_{2},\ldots,e_{m}\}$, where $m$ represents the total number of edits. Edits occurring in $m$ consecutive positions starting from the first hop are classified as pre (where $1\leq m\leq n$), while those occurring from the $(n-m+1)^{\text{th}}$ to the $n^{\text{th}}$ position are labeled post (also with $1\leq m\leq n$). Edits not including the first and last hops are categorized as mid.
@@ Appendix D Addition Experimental Settings
D.2 Experimental Settings
@@
Appendix E Additional Experimental Results
Model Method Average Accuracy 2-hop 3-hop 4-hop
GPT-J-6B Base 42.83 48.9 30.7 48.9
FT 1.9 3.7 1.4 0.5
MEND 11.5 13.9 11.3 9.5
ROME 18.1 33.8 9.1 11.4
MEMIT 12.3 22.5 6.0 8.4
PMET 17.04 26.65 12.76 11.7
IFMET (ours) 31.01 44.06 23.58 25.4
Table 5: Performance comparing the baselineand our method with CoT on multi-hop questions in MQuAKE-CF-3k, categorized by hop counts of 2, 3, 4.
Model Method Average Accuracy 1-Edit 2-Edit 3-Edit 4-Edit
GPT-J-6B Base 42.83 36.96 45.27 46.85 48.51
FT 1.9 4.2 0.7 0.3 0.0
MEND 11.5 16.0 11.0 7.3 4.4
ROME 18.1 23.8 20.9 9.0 2.6
MEMIT 12.3 20.5 9.8 5.5 2.6
PMET 17.04 22.63 16.74 11.19 7.84
IFMET (ours) 31.01 30.26 35.21 24.30 31.72
Table 6: Performance comparing the baselineand our method with CoT on multi-hop questions in MQuAKE-CF-3k, categorized by the number of edits 1, 2, 3, 4.
Figure 6: Comparison of different methods across batch sizes, hop numbers, and edit instances.
SELECT ?subject ?subjectLabel ?predicate ?predicateLabel WHERE
?subject ?predicate wd:ss.
FILTER (?predicate IN (wdt:relation))
SERVICE wikibase:label bd:serviceParam wikibase:language "en".
LIMIT 50
Table 7: The template of the SPARQL Query we used for the supplementary triplets.

### v3

@@ 17       $W^{l}\leftarrow W^{l}+\Delta^{l}$
where $K_{0}=\left[k^{1}_{0}\,|\,k^{2}_{0}\,|\,\cdots\,|\,k^{N}_{0}\right]$ and $V_{0}=W^{out}K_{0}$ contain all the knowledge we want to preserve, $K_{E}=\left[k^{1}_{e}\,|\,k^{2}_{e}\,|\,\cdots\,|\,k^{E}_{e}\right]$ is the matrix containing the edits we try to make and $V_{e}=[v^{*}_{e_{1}}\ |\dots\ |\ v^{*}_{e_{E}}]$ represents the target representations of the new knowledge. ($K_{E},V_{E}$) corresponds to the edited fact set $\{\left(s_{i},r_{i},o^{*}_{i}\right)|i=1,2,\cdots,E\}$. We consider the target weight $\hat{W}$ as the sum of the original weight $W^{out}$ and the incremental weight $\Delta$, as explicated in (Li et al., 2024c), a closed-form solution to the incremental weight can be derived:
@@
The primary differences between the first and furtherance edits are reflected in the edit prompt $T_{e}$ and the layers edited. For example, for the edit instance $e=(s,r,o\to o*)$, the first edit utilized a single-hop edit template $T_{e}=T_{r}(s)$ provided by MQuAKE to edit layers [3,8] of the GPT-J model in the subject last token position. For the furtherance edit, a two-hop prompt $T_{e}=T_{C}(s^{\prime})$ composed of a support case $(s^{\prime},r,s)$ and $(s,r,o^{*})$, and this two-hop edit prompt was applied to edit layers [16,20] of the GPT-J model in the last token position.
Appendix E Addition Experimental Settings
E.1 Criteria for classifying dataset into pre, mid, and post.
Consider a multi-hop question composed of $n$ triples. We define the positions of edits (with the index starting from 1) as the set $\{e_{1},e_{2},\ldots,e_{m}\}$, where $m$ represents the total number of edits. Edits occurring in $m$ consecutive positions starting from the first hop are classified as pre (where $1\leq m\leq n$), while those occurring from the $(n-m+1)^{\text{th}}$ to the $n^{\text{th}}$ position are labeled post (also with $1\leq m\leq n$). Edits not including the first and last hops are categorized as mid.
@@ Appendix E Addition Experimental Settings
E.2 Experimental Settings
When constructing the multi-hop edit prompts, for each edit case, no more than three triplets per relation were added. The relation types of the multi-hop edit prompts set are the same as MQuAKE. We set the edit batch sizes to 1.
@@ E.2 Experimental Settings
Appendix F Ablation Study
Method Stage Edit Prompt Position Layers
Previous Only one Single-hop Subject Last Token Shallow
IFMET
First
Furtherance
Single-hop
Multi-hop
Subject Last Token
Last Token
Shallow
Deeper
Table 6: The main difference between IFMET and previous methods. The term Stage refers to the phases of the editing process, Data denotes the query utilized for editing, Position specifies the token position where the editing is applied, and Layers indicate the edited layers.
Edits Editor Multi-hop Efficacy Specificity Paraphrase
1 IFMET 28.38 ($\uparrow$78.0%) 99.56 ($\uparrow$12.8%) 69.87 ($\downarrow$11.1%) 90.17 ($\uparrow$5.3%)
w/o $First$ 23.14 ($\uparrow$45.1%) 66.59 ($\downarrow$24.6%) 62.61 ($\downarrow$20.3%) 41.48 ($\downarrow$51.6%)
w/o $Multi$ 17.69 ($\uparrow$10.9%) 100.00 ($\uparrow$13.3%) 77.71 ($\downarrow$1.2%) 86.24 ($\uparrow$0.7%)
w/o $Last$ 18.12 ($\uparrow$13.6%) 88.21 ($\uparrow$0.0%) 78.60 ($\downarrow$0.0%) 86.24 ($\uparrow$0.7%)
w/o $Deeper$ 15.07 ($\downarrow$5.4%) 99.56 ($\uparrow$12.8%) 70.31 ($\downarrow$10.6%) 86.90 ($\uparrow$1.5%)
PMET 15.94 88.21 78.60 85.59
100 IFMET 27.07 ($\uparrow$64.8%) 96.29 ($\uparrow$8.1%) 69.89 ($\downarrow$9.4%) 84.28 ($\uparrow$3.8%)
w/o $First$ 22.71 ($\uparrow$35.1%) 73.36 ($\downarrow$17.8%) 69.21 ($\downarrow$10.2%) 34.72 ($\downarrow$57.3%)
w/o $Multi$ 17.25 ($\uparrow$2.6%) 99.13 ($\uparrow$11.3%) 76.63 ($\uparrow$0.6%) 84.06 ($\uparrow$3.5%)
w/o $Last$ 15.94 ($\downarrow$5.2%) 89.08 ($\downarrow$0.0%) 76.85 ($\uparrow$0.3%) 81.55 ($\uparrow$0.4%)
w/o $Deeper$ 16.16 ($\downarrow$3.9%) 99.56 ($\uparrow$11.8%) 74.67 ($\downarrow$3.2%) 81.00 ($\downarrow$0.3%)
PMET 16.81 89.08 77.07 81.22
Table 7: The results of the ablation experiments on GPT-J-6B model using a subset of MQuAKE-3K. Both the percentages of decrease($\downarrow$) and increase($\uparrow$) are calculated relative to PMET as the baseline. The most significant performance decline is highlighted in red and the most significant performance increase is highlighted in green.
Our mechanism analysis has identified that the existing editing methods with only single-hop edit prompts fail to adequately modify knowledge in the deeper MLP layers, resulting in poor performance on multi-hop factual recall tasks. Additionally, our findings suggest that implicit multi-hop step dependencies rely on the knowledge provided by these deeper MLP layers. Based on these interpretability results related to the mechanistic difference, we propose the IFMET. Given the distinctions between IFMET and other existing methods, we highlight four key components, especially in the furtherance edit: two-stage modification, the use of multi-hop edit prompt, editing the last token position, and updating knowledge in the deeper-layer MLPs during the second stage, as illustrated in the table 6.
To investigate whether the IFMET method effectively balances the requirements of general knowledge editing and multi-hop fact recall tasks, we constructed the paraphrase set and neighborhood set for a subset of the MQuAKE-3K dataset , following the approach used in the COUNTERFACT dataset (Meng et al., 2022b). We conducted experiments under two configurations: edit batch = 1 and edit batch = 100 and focused on analyzing the roles of these key components on the following metrics:
  • Efficacy measures whether an edit has been successfully applied to a model and whether the model can successfully answer the single-hop fact recall prompt. It is calculated as the percentage of edits where $P(\text{edited answer})>P(\text{unedited answer})$ for a given query prompt used during model editing.
  • Multi-hop measures whether the multi-hop question related to the edited knowledge can be successfully answered by the model.
  • Paraphrase evaluates the model’s generalization ability under an edit. It is defined as the percentage of edits where $P(\text{edited answer})>P(\text{unedited answer})$ for paraphrases of the query prompt.
  • Specificity assesses the locality of the model editing, i.e., whether the edit of a specific fact affects other facts stored within the model. Specificity score is defined as the percentage of facts in the neighborhood of the edited fact that remain unchanged after the edit.
In Table 7, we have utilized the PMET as a baseline to assess method performance. The importance of each component is reflected through comparisons of performance improvements over PMET. PMET’s performance exemplifies a single-stage locate-then-edit approach editing shallow MLPs based on a single-hop prompt. From the analysis of the ablation experiments, we derive the following conclusions:
  • IFMET: Firstly, it can be observed that the implementation of IFMET achieves the best performance in Multi-hop Acc. In the second stage of editing, we employ multi-hop edit prompts alongside deep MLP editing techniques. Across all the experimental tables mentioned, IFMET consistently demonstrates a substantial improvement in inferential performance compared to PMET. As the edit batch size increases, IFMET continues to sustain performance improvements. However, it is worth noting that the IFMET framework also leads to a slight decrease in the specificity metric, indicating that more extensive modifications to the model’s knowledge can somewhat compromise specificity. But, as mentioned in previous works, the pursuit of a balanced performance across the four metrics—Multi-hop Efficacy, Specificity, Paraphrase—we believe IFMET performs exceptionally well in this regard as well.
  • w/o $First$: When only applying the second edit stage in IFMET, which modifies the deeper layers using a multi-hop edit prompt, it effectively enhances performance on multi-hop reasoning tasks. However, the absence of first-stage editing results in unchanged knowledge in the earlier layers, leading to poor performance in single-hop fact recall tasks. This highlights the importance of the two stages, where each stage is indispensable.
  • w/o $Last$ demonstrated the importance of editing the last token position in the second stage. This is consistent with the findings from the mechanism exploration in Section 3.1, where subject information converges at the last token position and has a causal effect on the retrieval of the knowledge about the final answer.
  • w/o $Multi$: This represents that, in the second editing stage, we continued to use the original single-hop edit prompt instead of the multi-hop edit prompt to edit the deeper MLP layers. The performance improvement in efficiency somewhat indicates that single-layer edit prompts can edit knowledge in deep MLP for single-hop scenarios, but there is no significant improvement in multi-hop performance. This suggests that, due to the difference in mechanisms, the knowledge used in implicit reasoning steps that have not been sufficiently modified in the deep MLP layers remains unedited. It also indicates that using single-hop prompts combined with deep MLP editing is ineffective, highlighting the critical importance of the multi-hop prompts.
  • w/o $Deeper$: In this setup, the second-stage editing was modified to use the multi-hop edit prompt combined with shallow MLP editing(rather than deeper MLP layers). If this also shows a significant performance improvement in multi-hop accuracy, it would indicate that merely expanding with the multi-hop edit prompt, without considering its mechanism on the deeper MLP layers, can enhance results. However, as observed across the table, there was a consistent minor fluctuation in performance (ranging from -5.4% to +3.9%). In contrast to IFMET’s own +70% improvement, this underscores the importance of editing the deeper MLP layers when using the multi-hop edit prompts set.
In light of the results from the ablation experiments w/o $Multi$ and w/o $Deeper$, we emphasize that it is essential to apply multi-hop edit prompts to the deeper MLP layers for knowledge editing, ensuring a mechanism match, in order to effectively improve the performance of multi-hop factual recall tasks. In conclusion, we can summarize that the four components proposed by IFMET are essential for performance improvement, also highlighting the importance and validity of our interpretability exploration work.
Appendix G Generalizability of IFMET
In this subsection, We explore the generalizability of our method from four key perspectives:
Figure 6: Causal Intervention result of MLP hidden state in last token position on LLaMA-2
Generalization to other models. We first extended the causal intervention experiments in Section 3 to the LLaMA-2-7B model. The results shown in Figure 6 demonstrate the consistency of interpretability analysis across models, demonstrating the critical role of deeper-layer MLPs in LLaMA-2 model for multi-hop fact recall tasks. Additionally, we repeated the ablation experiments on LLaMA-2-7B to evaluate the generalizability of IFMET. The results shown in Table 8 are consistent with those observed on GPT-J, highlighting the importance of the four components in IFMET as well as the superiority of the method itself on LLaMA-2-7B model. Considering both the interpretability analysis and experimental outcomes, we conclude that our analysis and method are equally applicable to larger models, such as LLaMA-2.
Edits Editor Multi-hop Efficacy Specificity Paraphrase
1 IFMET 28.38 ($\uparrow$73.3%) 99.78 ($\uparrow$13.7%) 65.50 ($\downarrow$10.8%) 75.00 ($\uparrow$23.1%)
w/o $First$ 25.76 ($\uparrow$57.3%) 56.55 ($\downarrow$35.6%) 62.18 ($\downarrow$15.3%) 39.08 ($\downarrow$35.9%)
w/o $Multi$ 19.43 ($\uparrow$18.6%) 98.68 ($\uparrow$12.4%) 70.35 ($\downarrow$4.2%) 69.00 ($\uparrow$13.2%)
w/o $Last$ 15.72 ($\downarrow$4.1%) 88.86 ($\uparrow$1.2%) 73.03 ($\downarrow$0.6%) 61.90 ($\uparrow$1.6%)
w/o $Deeper$ 15.28 ($\downarrow$6.8%) 96.72 ($\uparrow$10.1%) 65.61 ($\downarrow$10.7%) 66.99 ($\uparrow$9.9%)
w $Multi_{model}$ 26.86 ($\uparrow$64.0%) 96.07 ($\uparrow$9.4%) 63.32 ($\downarrow$13.8%) 74.24 ($\uparrow$21.8%)
PMET 16.38 87.77 73.41 60.92
100 IFMET 27.29 ($\uparrow$76.0%) 97.82 ($\uparrow$4.2%) 65.50 ($\downarrow$9.9%) 84.17 ($\uparrow$14.2%)
w/o $First$ 24.67 ($\uparrow$59.2%) 64.41 ($\downarrow$31.4%) 63.97 ($\downarrow$12.0%) 41.81 ($\downarrow$43.3%)
w/o $Multi$ 15.07 ($\downarrow$2.8%) 99.34 ($\uparrow$5.8%) 71.83 ($\downarrow$1.2%) 76.64 ($\uparrow$4.0%)
w/o $Last$ 13.97 ($\downarrow$9.9%) 94.32 ($\uparrow$0.4%) 72.14 ($\downarrow$0.8%) 74.67 ($\uparrow$1.3%)
w/o $Deeper$ 15.94 ($\uparrow$2.8%) 96.51 ($\uparrow$2.8%) 69.48 ($\downarrow$4.4%) 75.44 ($\uparrow$2.3%)
w $Multi_{model}$ 22.49 ($\uparrow$45.1%) 96.07 ($\uparrow$2.3%) 63.32 ($\downarrow$12.9%) 79.26 ($\uparrow$7.5%)
PMET 15.50 93.89 72.66 73.69
Table 8: The results of the ablation experiments on LLaMA-2-7B model using a subset of MQuAKE-3K. w $Multi_{model}$ represents the multi-hop edit prompts generated by model itself to modify deeper MLPs. Both the percentages of decrease($\downarrow$ ) and increase($\uparrow$) are calculated relative to IFMET as the baseline.
Generalization to other edit batch number. In addition to editing a single instance at a time, we also tested scenarios where 1,000 and 3,000 instances are edited together. The results, shown in Figure 7, demonstrate that IFMET outperforms existing methods across all edit batch sizes.
Construction of the multi-hop prompts. In IFMET, we emphasize the importance of constructing a multi-hop edit prompt for each edit instance. It is important to note that any other valid knowledge base can replace WikiData and SPARQL. A straightforward alternative is to treat the model itself as a reliable knowledge base for extracting relevant knowledge.
The results of substituting the multi-hop edit prompt with one generated by LLaMA-2 itself for editing are also shown in Table 8 called w $Multi_{model}$. The results show a significant improvement over the one-stage PMET, with performance trends aligning closely with those of IFMET. Notably, minimal effort was invested in designing the knowledge retrieval prompt, and no additional filtering or preprocessing was applied. This suggests that the multi-hop edit prompts generated by the model represent a relatively low-quality version, effectively serving as a lower bound for the method’s performance across various metrics. Despite this, it still outperforms existing one-stage methods. This highlights the inherent superiority of the IFMET framework and demonstrates the feasibility of using the model itself to construct the prompts set.
Model Method Time
GPT-J-6B MEMIT 4.5s
PMET 5.0s
IFMET 9.7s
LLaMA-2-7B MEMIT 2.1s
PMET 2.0s
IFMET 3.4s
Table 9: The average time required to edit a single case varies across methods. For the two one-stage methods, MEMIT and PMET, this corresponds to the process of optimizing the shallow-layer MLPs using single-hop queries. For IFMET, the process includes updating the deeper-layer MLPs using two-hop edit prompts.
Time complexity of IFMET. We compared the time complexity of IFMET with that of the one-stage PMET method it builds upon, the result is shown in Table 9. On average, the time required to perform a complete edit for a single case on GPT-J using IFMET was approximately 2.5$\times$ that of PMET. For LLaMA-2, the time required was about 1.5$\times$ that of PMET. We believe this is within an acceptable range, and as the editing speed of the single-stage method improves, the IFMET framework will correspondingly become faster.
G.1 Comparison with Weight-Preserving Methods
Although there have been some weight-preserving editing methods(e.g. RAG-based Methods) accessing good performance for multi-hop question answering in the KE scenario, we believe that exploring the locate-then-edit methodology remains meaningful for several reasons:
  1. 1.
    From the perspective of understanding internal knowledge utilization: The mechanisms underlying a model’s use of internal knowledge differ fundamentally from those governing the use of external knowledge (Jin et al., 2024). Investigating the potential of locate-then-edit methods holds significant value for advancing the interpretability of internal knowledge processes, laying the groundwork for deeper insights and practical implementations. Additionally, we believe this approach enables a more fundamental and precise modification of knowledge.
  2. 2.
    From a practical standpoint: Methods based on retrieval-augmented generation (RAG) require providing extensive contextual input tokens, posing substantial challenges in terms of computational efficiency and hardware demands. And these methods face several challenges. Instead of injecting knowledge into LLMs, they retrieve related facts stored in memory for editing. As a result, their retrieval success rates become crucial, particularly when managing complex real-world scenarios involving exponential growth in knowledge updates. Moreover, we argue that an over-reliance on modifying knowledge through external contexts introduces security risks, as it may be exploited for data theft and attacks (Upadhayay et al., 2024), especially in real-world applications.
Appendix H Additional Experimental Results
Figure 7: Multi-hop Acc Performance Comparison of different methods across edit batch sizes on MQUAKE-3K. Base in this table represents unmodified GPT-J-6B model, and we report its performance on unedited answer.
(a) Two-hop
(b) Single-hop
Figure 8: LogitLens results of the last token position at different layers. (a) Yellow line represents the information containing implicit subject $s_{2}$, i.e., $\text{Info}(h_{l},s_{2})$. Blue line represents the information for the final answer, i.e., $\text{Info}(h_{l},o_{2})$. (b) Yellow line represents the information of subject $s$, i.e., $\text{Info}(h_{l},s)$, and Blue line represents the information of the answer $o$, i.e., $\text{Info}(h_{l},o)$. Larger versions of the sub-figures are available in the Appendix.
Appendix I Prompts And Templates
@@
(In-context-learning examples)
Input: The country that has nationals <mask> is located in the continent of Asia
Output: Hitomi Yaida
Input: The country that has nationals <mask> has the official language of Italian
Output: Giorgio Chiellini
Input: The university where <mask> was educated located its headquarters in the city of Vienna
Output: Michael Haneke
Input: The country that has nationals <mask>, its capital is Washington
Output: Lou Pearlman
Input: The person who found <mask> is a citizen of United States of America
Outout: Microsoft
Input: The creator of <mask> hails from Italy
Output: Ferrari
Input: The author of <mask> is a citizen of United States of America
Output: Holly Potter
Input: The person who discovered <mask> lives in Germany
Output: Volkswagen
Input: question
Table 13: The template of the prompt we used for asking LLaMA-2-7B to generate the multi-hop edit prompt
SELECT ?subject ?subjectLabel ?predicate ?predicateLabel WHERE
?subject ?predicate wd:ss.
FILTER (?predicate IN (wdt:relation))
SERVICE wikibase:label bd:serviceParam wikibase:language "en".
LIMIT 50
Table 14: The template of the SPARQL Query we used for prefix fact recall triplets for a specific subject.